# Polarity transitions induced by symmetry-breaking outer boundary heat flux in rapidly rotating dynamos

### v1

@@ 3.1 The effect of equatorially symmetric and anti-symmetric heat flux patterns on polarity transitions
30 645.16 176 176 782 0.07 25 38 8 4.18 4.5 2 44.72 5.69 0.48 62 0.77 D
Table 1: Summary of the main input and output parameters of a few dynamo simulations. Here, $Ra_{V}$ is the modified vertical Rayleigh number, $Ra_{V,crt}$ is the modified vertical critical Rayleigh number for onset of nonmagnetic convection, $q^{\star}$ is a dimensionless measure of the boundary heterogeneity defined in (42), $N_{r}$ is the number of radial grid points, $l_{\max}$ is the maximum spherical harmonic degree, $Rm$ is the magnetic Reynolds number, $Ro_{\ell}$ is the local Rossby number, $l_{C}$ and $l_{E}$ are the mean spherical harmonic degrees of convection and energy injection respectively, $\overline{m}$ is the mean spherical harmonic order in the range $l\leq l_{E}$, $\bar{k}_{s}$ and $\bar{k}_{z}$ are the mean $s$ and $z$ wavenumbers in the range $l\leq l_{E}$, $E_{k}$ and $E_{m}$ are the time-averaged total kinetic and magnetic energies, $f_{dip}$ is the relative dipole field strength, $Ra_{\ell,H}/Ra_{\ell}$ is the ratio of the local horizontal to the resultant Rayleigh numbers (obtained from (51) and (52)), and $B^{2}_{rms}$ is the measured mean square value of the field in the spherical shell. Types ‘D’, ‘R’, and ‘M’ denote dipolar, reversing and multipolar dynamo states, respectively. The other dynamo parameters are $E=1.2\times 10^{-5}$, $Pm=Pr=1$.
$Ra_{V}$ $Ra_{V}/Ra_{V,crt}$ $q^{\star}$ $N_{r}$ $l_{max}$ $Rm$ $Ro_{\ell}$ $l_{C}$ $l_{E}$ $\overline{m}$ $\bar{k}_{s}$ $\bar{k}_{z}$ $E_{k}$ $E_{m}$ $f_{dip}$ $B^{2}_{rms}$ $Ra_{\ell}$ $\varLambda$ $Ra_{\ell,H}/Ra_{\ell}$
×$\times 10^{5}$$\times 10^{5}$$\times 10^{4}$
@@
Table 2: Summary of the main input and output parameters in the reversing dynamo simulations considered in this study for $Y_{2}^{1}$ heat flux condition. Here, $Ra_{V}$ is the modified Rayleigh number, $Ra_{V,crt}$ is the modified vertical critical Rayleigh number for onset of nonmagnetic convection, $q^{\star}$ is a dimensionless measure of the boundary heterogeneity defined in (42), $N_{r}$ is the number of radial grid points, $l_{max}$ is the maximum spherical harmonic degree, $Rm$ is the magnetic Reynolds number, $Ro_{\ell}$ is the local Rossby number, $l_{C}$ and $l_{E}$ are the mean spherical harmonic degrees of convection and energy injection respectively (defined in (43)), $\overline{m}$ is the mean spherical harmonic order in the range $l\leq l_{E}$, $\bar{k}_{s}$ and $\bar{k}_{z}$ are the mean $s$ and $z$ wavenumbers in the range $l\leq l_{E}$, $E_{k}$ and $E_{m}$ are the time-averaged total kinetic and magnetic energies defined in (44), $f_{dip}$ is the relative dipole field strength, $Ra_{\ell}$ is the local Rayleigh number defined in (51) and $\varLambda$ is the peak Elsasser number obtained from the square of the measured peak field when slow MAC waves cease to exist in the run at high $q^{\star}$.
Since the ratio of anti-symmetric to total kinetic energy remains constant even at high $q^{\star}$ (figure 8), it is apparent that equatorially anti-symmetric conditions induce polarity transitions not by breaking the symmetry of the fluid columns. Before analysing the the role of wave motions in polarity transitions, we examine the mean temperature and velocity fields produced by the heterogeneous boundary heat flux, which give useful comparisons with the linear convection model.
Figure 8: Ratio of anti-symmetric to total kinetic energy for $q^{\star}=0$ (black), $q^{\star}=15$ (red), $q^{\star}=18$ (blue) and $q^{\star}=20$ (green) with the $Y_{2}^{1}$ heat flux heterogeneity at the outer boundary. The other dynamo parameters are $Ra=2500,E=1.2\times 10^{-5},Pm=Pr=1$.
3.2 Mean temperature and velocity fields
In line with that obtained in equation (33) in the Cartesian model, the resultant mean temperature gradient is given by
@@ 3.2 Mean temperature and velocity fields
where $\beta_{s}=\partial T_{0}/\partial s<0$ in an unstably stratified layer with homogeneous boundary heat flux (figure 9(a)) and $\beta_{z}=\partial T_{0}/\partial z$ takes both positive and negative values for an equatorially anti-symmetric $Y_{2}^{1}$ heat flux heterogeneity at the outer boundary (figure 9(b)). The temperature gradients are calculated from a steady state just below nonmagnetic convective onset ($Ra_{V}=30$, whereas onset occurs at $Ra_{V}=31$ for $q^{\star}=18$) so that the perturbations do not affect the calculation of the mean gradients. To calculate the resultant gradient $\beta$, the average $s$ and $z$ wavenumbers as defined in (53) are used (see §3.3 below for the procedure). The resultant $\beta$ calculated at $Ra_{V}=30$ and the resultant $\beta$ averaged over $\approx$ 0.5 diffusion time from the strongly supercritical state at $Ra_{V}=2500$ exhibit similar structures as well as nearly equal maximum and minimum values (see figure S1 in the Supplementary Material). In effect, regions of $\beta>0$ are stably stratified, consistent with the suppression of convection even at $Ra_{V}=2500$ (figures 9 (d) and (f)). On the other hand, regions of $\beta<0$ support convection. Interestingly, $\beta_{\phi}=\partial T_{0}/\partial\phi$ (figure 9(c)) has practically no influence on $\beta$, which justifies our original assumption of ignoring variations with respect to the non-preferred horizontal direction (see (14), §2.2).
@@
The convective region produced by $\beta_{s}<0$ and $\beta_{z}>0$ with the equatorially anti-symmetric heat flux at the boundary corresponds to the unstably stratified system modelled by the equatorial radial configuration of the Cartesian linear model in §2.1. An averaged $\beta_{z}$, obtained by taking the average of the positively signed $\partial T_{0}/\partial z$ within the spherical shell, is proportional to the intensity of the anti-symmetric heterogeneity. Likewise, the averaged $\beta_{s}$ is obtained by taking the average of $\partial T_{0}/\partial s$ within the shell. The averaged resultant gradient $\beta$ is then evaluated from (46) using the $s$ and $z$ wavenumbers. Figure 10(a) shows the azimuthal variation of the mean temperature gradients at cylindrical radius $s=1$ and at the section $z=0.4$ below the equator for the $Y_{2}^{1}$ heterogeneity with $q^{\star}=18$. Figure 10(b) gives the meridional plot of the steady mean flow $u_{0}$ averaged over the unstably stratified longitudes in figure 10(a) ($\phi=1.31\,\pi$ to $0.31\,\pi$ through $\phi=0$). In figure 10(c), the magnitude of this equatorially anti-symmetric flow measured at $s=1$ shows a fair agreement with the theoretical value of the thermal wind calculated from equation (3) in dimensionless units,
@@
Figure 9: Horizontal section plots at $z=0.4$ below the equator showing (a) $\beta_{s}$ for homogeneous boundary heating, (b) $\beta_{z}$ for $q^{\star}=18$, (c) $\beta_{\phi}$ for $q^{\star}=18$, and (d) resultant gradient $\beta$ for $q^{\star}=18$. The temperature gradients are calculated at vertical Rayleigh number $Ra_{V}=30$, which is a state near the onset of nonmagnetic convection. A $Y_{2}^{1}$ heterogeneity in outer boundary heat flux is applied. The orange-green coloured strip at the periphery of plot (b) represents $\partial T_{0}/\partial r$ at the outer boundary. Panels (e) and (f) show snapshots of the axial velocity $u_{z}$ at $q^{\star}=0$ and $q^{\star}=18$ respectively for dynamo simulations at $Ra_{V}=2500$. The other parameters are $E=1.2\times 10^{-5}$, and $Pm=Pr=1$.
@@
Figure 10: (a) Values of $\beta_{s}$ (blue), $\beta_{z}$ (red) and the resultant $\beta$ (black) at cylindrical radius $s=1$ at the section $z=0.4$ below the equator for $q^{\star}=18$. (b) Meridional sections plot of $u_{0}$ for $q^{\star}=18$. The plot shows $\phi$-averaged values of the mean flow $u_{0}$ in the unstably stratified region where $\beta<0$ ($\phi=1.31\,\pi$ to $0.31\,\pi$ through $\phi=0$) in (a). (c) The values of $u_{0}$ along a vertical line passing through cylindrical radius $s=1$ (circles) are compared with the theoretical dimensionless thermal wind in equation (47). The parameters are $Ra_{V}=30$, $E=1.2\times 10^{-5}$, and $Pm=Pr=1$.
@@ 3.3 The role of slow MAC waves in the dipole–multipole transition
$\omega_{C}^{2}=\frac{Pm^{2}}{E^{2}}\frac{k_{z}^{2}}{k^{2}},\quad\omega_{M}^{2}=\frac{Pm}{E}(\bm{B}\cdot{\bm{k}})^{2},\quad-\omega_{A,V}^{2}=\frac{Pm^{2}Ra_{V}}{Pr\,E}\,\frac{{k_{h}}^{2}}{k^{2}},\quad\omega_{A,H}^{2}=\frac{Pm^{2}Ra_{H}}{2Pr\,E}\,\frac{{k_{h}}^{2}}{k^{2}},\quad$ (49)
where $k_{s}$, $k_{\phi}$ and $k_{z}$ are the radial, azimuthal and axial wavenumbers in cylindrical coordinates $(s,\phi,z)$, $k_{\phi}=m/s$, where $m$ is the spherical harmonic order, $k^{2}=k_{s}^{2}+k_{\phi}^{2}+k_{z}^{2}$, and $k_{h}$ is the horizontal wavenumber in the equatorial region defined by $k_{h}^{2}=k_{\phi}^{2}+k_{z}^{2}$. The magnetic (Alfvén) wave frequency $\omega_{M}$ is based on the three components of the measured magnetic field at the peak-field location. The wavenumber $k_{\phi}$ is evaluated at $s=1$, approximately mid-radius of the spherical shell. The horizontal Rayleigh number is given by
@@
Figure 11: Variation of the squares of the fundamental frequencies with $q^{\star}$ and $Ra_{\ell,H}/Ra_{\ell}$ (within brackets) for (a) $Ra_{V}=2500$, $Y_{2}^{1}$ heat flux heterogeneity, and (b) $Ra_{V}=2500$, $Y_{2}^{2}$ heat flux heterogeneity. The dotted vertical line marks the polarity-reversing state that lies between the dipolar and multipolar regimes. The other dynamo parameters are $E=1.2\times 10^{-5}$, $Pm=Pr=1$.
In figure 11, the magnitudes of fundamental frequencies are shown as a function of $q^{\star}$ and $Ra_{\ell,H}/Ra_{\ell}$ for heterogeneous heat flux boundary conditions. The frequencies and local Rayleigh numbers are computed from (49), (51) and (52) using the mean values of the $s$, $\phi$ and $z$ wavenumbers in the saturated state of the dynamo runs in the energy-containing scales ($l\leq l_{E}$). For example, real space integration over $(s,\phi)$ gives the kinetic energy as a function of $z$, the Fourier transform of which gives the one-dimensional spectrum $\hat{u}^{2}(k_{z})$. Subsequently, we obtain,
$\bar{k}_{z}=\frac{\Sigma k_{z}\,\hat{u}^{2}(k_{z})}{\Sigma\hat{u}^{2}(k_{z})}.$ (53)
A similar approach yields $\bar{k}_{s}$ and $\overline{m}$.
@@
In Figure 11(a), we begin with a strong field dynamo and increase the $Y_{2}^{1}$ heat flux heterogeneity, which is reflected in the horizontal buoyancy frequency $\omega_{A,H}$, while keeping the vertical buoyancy constant at $Ra_{V}=2500$. The slow MAC waves are present at relatively low $q^{\star}$ when $|\omega_{M}|>|\omega_{A}|$. At $q^{\star}=18$ ($Ra_{\ell,{H}}/Ra_{\ell}\approx 0.62$), the magnitude of the resultant buoyancy frequency $|\omega_{A}|$ approximately matches that of the Alfvén frequency $|\omega_{M}|$, causing the slow wave frequency $\omega_{s}$ to go to zero, which in turn results in a transition to the multipolar state through reversals. In contrast, the equatorially symmetric $Y_{2}^{2}$ heat flux boundary condition has zero horizontal buoyancy at the equator, and even at large $q^{\star}$, the condition $|\omega_{A}|>|\omega_{M}|$ is never met (figure 11(b)).
@@
Figure 12: Square of the absolute values of wave frequencies plotted from the saturated state of the dynamo run for $q^{\star}$= 0 (a), 10 (b) and 18 (c). The shaded grey area shows the range of scales where the helicity of the dynamo run is greater than that of the equivalent non-magnetic run. The other dynamo parameters are $Ra_{V}=2500$, $E=1.2\times 10^{-5}$, $Pm=Pr=1$.
@@
The top panel of figure 12 displays the square of the fundamental frequencies for a range of spherical harmonic order $m$ for (a) $q^{\star}=0$, (b) $q^{\star}=10$, and (c) $q^{\star}=18$ in the saturated state. Dynamo simulations with $q^{\star}=0$ and $q^{\star}=10$ exhibit stable axial dipoles while the simulation with $q^{\star}=18$ gives a reversing dynamo. The shaded area indicates the range of $m$ where the kinetic helicity in the dynamo exceeds that of the equivalent non-magnetic run. (For the governing equations of equivalent non-magnetic run, see Majumder et al., 2024). The black line represents the square of the slow MAC wave frequency under the condition $|\omega_{C}|>|\omega_{M}|>|\omega_{A}|$. As $q^{\star}$ increases, the range of $m$ satisfying this condition narrows, and for $q^{\star}=18$, this condition is not satisfied for any $m$. The bottom panel of figure 12 gives the spectral distribution of the power supplied to the poloidal component of the axial dipole field $B_{10}^{P}$ (e.g. Buffett & Bloxham, 2002) for the three $q^{\star}$ values,
[…]
@@
where $\bm{u}$ and $\bm{B}$ have the same $m$ (Bullard & Gellman, 1954). The dipole power peaks at $q^{\star}=0$ and $10$ within the range of $m$ where slow MAC waves are generated and the helicity of the dynamo is greater than that of the equivalent non-magnetic simulation. In contrast, the run with $q^{\star}=18$ shows no such peak since there is no helicity generation at any $m$. The $f_{dip}$ values for $Ra=2500$ decrease with increasing heterogeneity, with a sharp drop observed at the polarity transition for $q^{\star}=18$ (table 1). These results indicate the crucial role of the slow MAC waves in helicity generation in the energy-containing scales, and in turn, dipole formation.

### v2

@@ 3.1 The effect of equatorially symmetric and anti-symmetric heat flux patterns on polarity transitions
30 645.16 176 176 782 0.07 25 38 8 4.18 4.52 31.44 35.61 0.54 156 3.18 0.77 D
Table 1: Summary of the main input and output parameters of a few dynamo simulations. Here, $Ra_{V}$ is the modified vertical Rayleigh number, $Ra_{V,\,c}$ is the modified vertical critical Rayleigh number for onset of nonmagnetic convection, $q^{*}$ is a dimensionless measure of the boundary heterogeneity defined in (45), $N_{r}$ is the number of radial grid points, $l_{\max}$ is the maximum spherical harmonic degree, RmRm is the magnetic Reynolds number, $Ro_{\ell}$ is the local Rossby number, $l_{C}$ and $l_{E}$ are the mean spherical harmonic degrees of convection and energy injection respectively, $\overline{m}$ is the mean spherical harmonic order in the range $l\leq l_{E}$, $\bar{k}_{s}$ and $\bar{k}_{z}$ are the mean $s$ and $z$ wavenumbers in the range $l\leq l_{E}$, $E_{k}$ and $E_{m}$ are the time-averaged total kinetic and magnetic energies, $f_{dip}$ is the relative dipole field strength, $Ra_{\ell,H}/Ra_{\ell}$ is the ratio of the local horizontal to the resultant Rayleigh numbers (obtained from (55) and (56)), and $B^{2}_{rms}$ is the measured mean square value of the field in the spherical shell. Types ‘D’, ‘R’, and ‘M’ denote dipolar, reversing and multipolar dynamo states, respectively. The other dynamo parameters are $E=1.2\times 10^{-5}$, $Pm=Pr=1$.
$Ra_{V}$ $\dfrac{Ra_{V}}{Ra_{V,\,c}}$ $q^{*}$ $N_{r}$ $l_{max}$ RmRm $Ro_{\ell}$ $l_{C}$ $l_{E}$ $\overline{m}$ $\bar{k}_{s}$ $\bar{k}_{z}$ $E_{k}$ $E_{m}$ $f_{dip}$ $B^{2}_{rms}$ $Ra_{\ell}$ $B^{2}_{peak}$ $\dfrac{Ra_{\ell,H}}{Ra_{\ell}}$
$\times 10^{5}$ $\times 10^{5}$ $\times 10^{4}$
@@
Table 2: Summary of the main input and output parameters in the reversing dynamo simulations considered in this study for $Y_{2}^{1}$ heat flux condition. These simulations have $|\omega_{A}/\omega_{M}|\geq 1$ throughout. Here, $Ra_{V}$ is the modified Rayleigh number, $Ra_{V,\,c}$ is the modified vertical critical Rayleigh number for onset of nonmagnetic convection, $q^{*}$ is a dimensionless measure of the boundary heterogeneity defined in (45), $N_{r}$ is the number of radial grid points, $l_{max}$ is the maximum spherical harmonic degree, RmRm is the magnetic Reynolds number, $Ro_{\ell}$ is the local Rossby number, $l_{C}$ and $l_{E}$ are the mean spherical harmonic degrees of convection and energy injection respectively (defined in (46)), $\overline{m}$ is the mean spherical harmonic order in the range $l\leq l_{E}$, $\bar{k}_{s}$ and $\bar{k}_{z}$ are the mean $s$ and $z$ wavenumbers in the range $l\leq l_{E}$, $E_{k}$ and $E_{m}$ are the time-averaged total kinetic and magnetic energies defined in (47), $f_{dip}$ is the relative dipole field strength, $Ra_{\ell}$ is the local Rayleigh number defined in (55) and $B^{2}_{peak}$ is the square of the measured peak field when slow MAC waves cease to exist when $|\omega_{A}|\approx|\omega_{M}|$.
The ratio of anti-symmetric to total kinetic energy remains nearly constant even at high $q^{*}$ (figure 8), so equatorially anti-symmetric boundary conditions do not induce polarity transitions by breaking the equatorial symmetry of the convection columns. Before analysing the role of wave motions in polarity transitions, we examine the mean temperature and velocity fields produced by the heterogeneous boundary heat flux, which give useful comparisons with the linear convection model.
Figure 8: Ratio of anti-symmetric to total kinetic energy for $q^{*}=0$ (black), $q^{*}=15$ (red), $q^{*}=18$ (blue) and $q^{*}=20$ (green) with the $Y_{2}^{1}$ heat flux heterogeneity at the outer boundary. The other dynamo parameters are $Ra=2500,E=1.2\times 10^{-5},Pm=Pr=1$.
3.2 Mean resultant temperature gradient and velocity field
In line with that obtained in equation (36) in the Cartesian model, the mean (superadiabatic) resultant temperature gradient is given by
@@ 3.2 Mean resultant temperature gradient and velocity field
where $\beta_{s}=\partial T_{0}/\partial s<0$ in an unstably stratified layer with homogeneous boundary heat flux (see figure 9(a)). In figure 9(b), $\beta_{z}=\partial T_{0}/\partial z$ is plotted for an equatorially antisymmetric $Y_{2}^{1}$ heat-flux heterogeneity at the outer boundary. The temperature gradients are calculated from a steady state just below nonmagnetic convective onset ($Ra_{V}=30$, whereas onset occurs at $Ra_{V}=31$ for $q^{*}=18$) so that the perturbations do not affect the calculation of the mean gradients. An averaged $\beta_{z}$, obtained by taking the average of the positively signed $\partial T_{0}/\partial z$ within the spherical shell, is proportional to the intensity of the anti-symmetric heterogeneity. Likewise, the averaged $\beta_{s}$ is obtained by taking the average of $\partial T_{0}/\partial s$ within the shell. The averaged resultant gradient $\beta$ is then evaluated from (49) using the average $s$ and $z$ wavenumbers defined in (53) (see figure 9(d)). The resultant $\beta$ calculated at $Ra_{V}=30$ and the resultant $\beta$ averaged over $\approx$ 0.5 diffusion time from the strongly supercritical state at $Ra_{V}=310$ exhibit similar structures as well as nearly equal maximum and minimum values (see figure S4 in the Supplementary Material). Convection is suppressed in stably stratified regions where $\partial T/\partial r>0$. In contrast, regions beneath $\partial T/\partial r<0$, where $\beta_{z}>0$ and $\beta<0$, support convection. This unstably stratified region is in focus in the analysis in §3.3 below – at moderately positive values of $\beta_{z}$, both fast and slow MAC waves are present, giving the dipolar dynamo regime; at large positive values of $\beta_{z}$, the slow MAC waves are selectively suppressed whereas the fast MAC waves persist, giving the reversing and multipolar dynamo regimes.
Equation (49) is derived from the theoretical model described in §2, which predicts the presence or absence of slow MAC waves subject to the frequency inequality $|\omega_{C}|>|\omega_{M}|,|\omega_{A}|>|\omega_{\eta}|$, with slow MAC waves present when $|\omega_{M}|>|\omega_{A}|$ and absent when $|\omega_{M}|\leq|\omega_{A}|$. Since the above frequency inequality does not predict the localized suppression of convection observed in figure 9(f), equation (49) can only be applied to regions where $\beta_{z}>0$. Interestingly, $\beta_{\phi}=\partial T_{0}/\partial\phi$ (figure 9(c)) has practically no influence on $\beta$, which justifies our original assumption of ignoring variations with respect to the non-preferred horizontal direction (see (15), §2.2). The convective region produced by $\beta_{s}<0$ and $\beta_{z}>0$ with the equatorially anti-symmetric heat flux at the boundary corresponds to the unstably stratified system modelled by the equatorial radial configuration of the Cartesian linear model in §2.1.
@@
Figure 10(a) shows the azimuthal variation of the mean temperature gradients at cylindrical radius $s=1$ and at the section $z=0.4$ below the equator for the $Y_{2}^{1}$ heterogeneity with $q^{*}=18$. The grey-shaded region indicates $\beta_{z}>0$, where the resultant $\beta$ is calculated from (49). Figure 10(b) gives the meridional plot of the steady mean flow $u_{0}$ ($\phi$-component of the velocity) averaged over the unstably stratified longitudes in figure 10(a) ($\phi=1.31\,\pi$ to $0.31\,\pi$ through $\phi=0$). In figure 10(c), the magnitude of this equatorially anti-symmetric flow measured at $s=1$ shows a fair agreement with the theoretical value of the thermal wind calculated from equation (4) in dimensionless units,
@@
Figure 9: Horizontal section plots at $z=0.4$ below the equator showing (a) $\beta_{s}$ for homogeneous boundary heating, (b) $\beta_{z}$ for $q^{*}=18$, (c) $\beta_{\phi}$ for $q^{*}=18$, and (d) resultant gradient $\beta$ for $q^{*}=18$. The temperature gradients are calculated at vertical Rayleigh number $Ra_{V}=30$, which is a state near the onset of nonmagnetic convection. A $Y_{2}^{1}$ heterogeneity in outer boundary heat flux is applied. The orange-green coloured strip at the periphery of plot (b) represents $\partial T_{0}/\partial r$ at the outer boundary. Panels (e) and (f) show snapshots of the axial velocity $u_{z}$ at $q^{*}=0$ and $q^{*}=18$ respectively for dynamo simulations at $Ra_{V}=2500$. The other parameters are $E=1.2\times 10^{-5}$, and $Pm=Pr=1$.
@@
Figure 10: (a) Values of $\beta_{s}$ (blue), $\beta_{z}$ (red) and the resultant $\beta$ (black) at cylindrical radius $s=1$ at the section $z=0.4$ below the equator for $q^{*}=18$. (b) Meridional sections plot of the $\phi$-component of the velocity, $u_{0}$ for $q^{*}=18$. The plot shows $\phi$-averaged values of $u_{0}$ in the unstably stratified region where $\beta<0$ ($\phi=1.31\,\pi$ to $0.31\,\pi$ through $\phi=0$) in the grey-shaded region of figure (a). (c) The values of $u_{0}$ along a vertical line passing through cylindrical radius $s=1$ (circles) are compared with the theoretical dimensionless thermal wind in equation (50). The parameters are $Ra_{V}=30$ (below onset), $E=1.2\times 10^{-5}$, and $Pm=Pr=1$. A $Y_{2}^{1}$ heat flux heterogeneity is applied at the outer boundary.
@@ 3.3 The role of slow MAC waves in the dipole–multipole transition
$\omega_{C}^{2}=\frac{Pm^{2}}{E^{2}}\frac{k_{z}^{2}}{k^{2}},\quad\omega_{M}^{2}=\frac{Pm}{E}(\bm{B}\cdot{\bm{k}})^{2},\quad-\omega_{A,V}^{2}=\frac{PmRa_{V}}{E}\,\frac{{k_{h}}^{2}}{k^{2}},\quad\omega_{A,H}^{2}=\frac{PmRa_{H}}{2\,E}\,\frac{{k_{h}}^{2}}{k^{2}},\quad$ (52)
where $k_{s}$, $k_{\phi}$ and $k_{z}$ are the radial, azimuthal and axial wavenumbers in cylindrical coordinates $(s,\phi,z)$, $k_{\phi}=m/s$, where $m$ is the spherical harmonic order, $k^{2}=k_{s}^{2}+k_{\phi}^{2}+k_{z}^{2}$, and $k_{h}$ is the horizontal wavenumber in the equatorial region defined by $k_{h}^{2}=k_{\phi}^{2}+k_{z}^{2}$. The $s$, $\phi$ and $z$ wavenumbers are calculated in the saturated state of the dynamo runs in the energy-containing scales ($l\leq l_{E}$). For example, real space integration over $(s,\phi)$ gives the kinetic energy as a function of $z$, the Fourier transform of which gives the one-dimensional spectrum $\hat{u}^{2}(k_{z})$. Subsequently, we obtain,
$\bar{k}_{z}=\frac{\Sigma k_{z}\,\hat{u}^{2}(k_{z})}{\Sigma\hat{u}^{2}(k_{z})}.$ (53)
A similar approach yields $\bar{k}_{s}$ and $\overline{m}$. The magnetic (Alfvén) wave frequency $\omega_{M}$ is based on the three components of the measured magnetic field at the peak-field location. The wavenumber $k_{\phi}$ is evaluated at $s=1$, approximately mid-radius of the spherical shell. The horizontal Rayleigh number is given by
@@
Figure 11: Variation of the squares of the fundamental frequencies with $q^{*}$ and $Ra_{\ell,H}/Ra_{\ell}$ (within brackets) for (a) $Ra_{V}=2500$, $Y_{2}^{1}$ heat flux heterogeneity, and (b) $Ra_{V}=2500$, $Y_{2}^{2}$ heat flux heterogeneity. The dotted vertical line marks the polarity-reversing state that lies between the dipolar and multipolar regimes. The other dynamo parameters are $E=1.2\times 10^{-5}$, $Pm=Pr=1$.
Figure 12: Variation of $f_{dip}$ with $|\omega_{A}/\omega_{M}|$ for different heat flux boundary conditions at two values of $Ra_{V}$. The vertical dotted line indicates $|\omega_{A}/\omega_{M}|=1$, while the horizontal dotted line marks $f_{dip}=0.35$, which has been proposed as the lower bound for the existence of dipole-dominated numerical dynamos (Christensen and Aubert, 2006). Types ‘D’, ‘R’, and ‘M’ denote dipolar, reversing and multipolar dynamo states, respectively. The other dynamo parameters are $E=1.2\times 10^{-5}$ and $Pm=Pr=1$.
@@
In figure 11, the magnitudes of fundamental frequencies are shown as a function of $q^{*}$ and $Ra_{\ell,H}/Ra_{\ell}$ for heterogeneous heat flux boundary conditions. The frequencies and local Rayleigh numbers are computed from (52), (55) and (56) using the mean values of the wavenumbers as shown in equation (53). In Figure 11(a), we begin with a strong field dynamo and increase the $Y_{2}^{1}$ heat flux heterogeneity, which is reflected in the horizontal buoyancy frequency $\omega_{A,H}$, while keeping the vertical buoyancy constant at $Ra_{V}=2500$. The slow MAC waves are present at relatively low $q^{*}$ when $|\omega_{M}|>|\omega_{A}|$. In the polarity-reversing state at $q^{*}=18$ ($Ra_{\ell,{H}}/Ra_{\ell}\approx 0.62$), the magnitude of the resultant buoyancy frequency $|\omega_{A}|$ approximately matches that of the Alfvén frequency $|\omega_{M}|$, causing the slow wave frequency $\omega_{s}$ to go to zero. While the generation of slow MAC waves ceases, the existing slow MAC waves are damped on the time scale $\delta^{2}/\eta$ (see §2.4), which is much shorter than the magnetic diffusion time $L^{2}/\eta$ because the ratio of the core depth to the length scale of buoyancy disturbances would be $L/\delta\sim 10^{2}$ (Sreenivasan and Maurya, 2021). Whether the rapid damping of the slow MAC waves causes the rapid decay of the axial dipole field during a polarity reversal (see figure 6 b) is at present an open question that requires further investigation. The equatorially symmetric $Y_{2}^{2}$ heat flux boundary condition has zero horizontal buoyancy at the equator, and even at large $q^{*}$, the condition $|\omega_{A}|>|\omega_{M}|$ is never met (figure 11(b)).
The variation of the parameter $f_{dip}$ against $|\omega_{A}/\omega_{M}|$ in figure 12 shows that reversing (R) and multipolar (M) states exist at values of $f_{dip}\leq 0.35$, the proposed lower bound for the existence of dipole-dominated numerical dynamos in the literature.
@@
Figure 13: Panels (a)–(c) (i) show the absolute values of wave frequencies plotted for the saturated state of the dynamo run. (a) $q^{*}=0$, (b) $q^{*}=10$ and (c) $q^{*}=18$ for the $Y_{2}^{1}$ heat flux heterogeneity at the outer boundary. The shaded grey area shows the range of scales where the helicity of the dynamo run is greater than that of the equivalent non-magnetic run. Panels (a)–(c) (ii) show the spectral distribution of the power supplied to the axial dipole, defined in (57). The other dynamo parameters are $Ra_{V}=2500$, $E=1.2\times 10^{-5}$, $Pm=Pr=1$.
@@
The top panel of figure 13 displays the square of the fundamental frequencies for a range of spherical harmonic order $m$ for (a) $q^{*}=0$, (b) $q^{*}=10$, and (c) $q^{*}=18$ in the saturated state for the equatorially anti-symmetric $Y_{2}^{1}$ heat flux pattern. Dynamo simulations with $q^{*}=0$ and $q^{*}=10$ exhibit stable axial dipoles while the simulation with $q^{*}=18$ gives a reversing dynamo. The shaded area indicates the range of $m$ where the kinetic helicity in the dynamo exceeds that of the equivalent non-magnetic run. (For the governing equations of equivalent non-magnetic run, see Majumder et al., 2024). The black line represents the square of the slow MAC wave frequency under the condition $|\omega_{C}|>|\omega_{M}|>|\omega_{A}|$. As $q^{*}$ increases, the range of $m$ satisfying this condition narrows, and for $q^{*}=18$, this condition is not satisfied for any $m$. The bottom panel of figure 13 gives the spectral distribution of the power supplied to the poloidal component of the axial dipole field $B_{10}^{P}$ (e.g. Buffett and Bloxham, 2002) for the three $q^{*}$ values,
$P_{\mathrm{10}}=\int_{V}\bm{B}^{P}_{\mathrm{10}}\cdot[\mathbf{\nabla}\times(\bm{u}\times\bm{B})_{m}]\,dV,$ (57)
@@
where $\bm{u}$ and $\bm{B}$ have the same $m$ (Bullard and Gellman, 1954). The dipole power peaks at $q^{*}=0$ and $q^{*}=10$ within the range of $m$ where slow MAC waves are generated and the helicity of the dynamo is greater than that of the equivalent non-magnetic simulation. In contrast, the run with $q^{*}=18$ shows no such peak since there is no helicity generation at any $m$. The $f_{dip}$ values for $Ra=2500$ decrease with increasing heterogeneity, with a sharp drop observed at the polarity transition for $q^{*}=18$ (table 1). These results indicate the crucial role of the slow MAC waves in helicity generation in the energy-containing scales, and in turn, dipole formation.